D owns all 30,000 shares of common stock of B. D has 60,000 shares of its own common stock outstanding. During the year D earns $200,000 none included from B, while B reports $150,000. Annual amortization of $10,000 is recognized each year on the consolidated worksheet based on acquisition date fair values. Both companies have convertible bonds outstanding. During the current year bond related interest expense (net of taxes) is $32,000 for D and 24,000 for B. D’s bonds can be converted into 10,000 shares. D owns none of B’s bonds.

Required:

Prepare a schedule which shows the earnings per share amounts that should D should report in its current consolidated income statement.
Answer #1

Since D owns 100% shares of B, we will directly add the entire income statement to D's consolidated income statement

Particulars D B Consolidated
Revenue (A) $        2,00,000 $        1,50,000 $       3,50,000
Amortization (B) $           10,000
EBIT (C = A - B) $       3,40,000
Interest (net of taxes) (D) $            32,000 $            24,000 $           56,000
Net income (E = C - D) $       2,84,000
No of shares (F) 60000 30000 90000
EPS (G = E/F) $               3.16

Now since D's bonds are convertible into 10,000 shares we need to include this in the calculation as Diluted EPS.

But before including we need to check whether the debt is dilutive or anti-dilutive.

We can check this by using the following formula, if the output is greater than the EPS calculated above, we do not calculated diluted EPS and take the basic EPS as the consolidated EPS

(convertible bond interest net of taxes/no. of convertible debt shares) =

= ($32,000/10,000) = $ 3.20 which is greater than 3.16 calculated above.

The bonds are anti-dilutive and hence we do not consider them in EPS calculation.